Acid rain over the Great Lakes has a pH of about 4.5 while acid rain over the West Coast has a pH of about 5.4. The lakes are then acidified by rainfall but are after treated through a neutralizing process called liming. This process involves adding limestone to the lake in large quantities.

a. What are the [H3O+] concentrations in each lake?
b. How many more times concentrated is the acid rain over the Great Lakes?

Please show all work/explain...I'd really like to understand how to get to the answer. THANKS
Answer #1

a)

pH = - log10 [H3O+] (where [H3O+] is the concentration of hydronium ions in Mol/Litre)

or, log10 [H3O+] = -pH

or, [H3O+] = 10(-pH)

  • For Great Lake, [H3O+] = 10(- 4.5) mol/litre

             Or, [H3O+] = 3.162 * 10-5 mol/litre

  • For West Coast, [H3O+] = 10(- 5.4) mol/litre

             Or, [H3O+] = 3.981 * 10-6 mol/litre

b)

The concentrations are in the order of 10-5 for Great Lake and 10-6 for West Coast (clearly, the factors 3.162 and 3.981 don’t have much saying in the numerical values of these orders).

Now, 10-6 * 10 = 10-5

Thus we can see, the acid rains over the Great Lake are almost 10 times more concentrated than that over the West Coast.

To be precise, it's (3.162 * 10-5 mol/litre) / (3.981 * 10-6 mol/litre) = .794 * 10 = 7.94 times more concentrated than the latter.